Jessusiskingforever:

The height of the tip of an airplane propeller above the ground once the airplane reaches full speed can be modelled by a sine function. At full speed, the propeller makes 200 revolutions per second. At t=0 the tip of the propeller is at its minimum height above the ground. Determine whether the instantaneous rate of change in height at t = 1/300 is a negative value, a positive value, or zero.

Narad:

\[\omega=2\pi*200=400\pi rad/s\] \[h(t)=\sin(400 \pi t-\frac{ \pi }{ 2 })\] \[h'(t)=400\pi \cos(400 pit-\frac{ \pi }{ 2 })\] when t=1/300, you calculate the value of h'(t) and you¡ll get your answer

Jessusiskingforever:

@narad wrote:
\[\omega=2\pi*200=400\pi rad/s\] \[h(t)=\sin(400 \pi t-\frac{ \pi }{ 2 })\] \[h'(t)=400\pi \cos(400 pit-\frac{ \pi }{ 2 })\] when t=1/300, you calculate the value of h'(t) and you¡ll get your answer
Thank you!